The preimeter of a rectangle is 200 inches. The length exceeds the width by 46 inches. Find the length and the width.
Answer by jojo14344(1513)

P= 2L + 2W ---------- eqn1
And L = W+46
So,
200=2(W+46) + 2w
200=2W+92+2W
4W=200-92
W=108/4
W=27 inches
For L = 27+46 = 73 inches
In doubt? Go back eqn1,
200= 2(73) + 2(27)
200=146+54
200=200
Thank you,
Jojo